MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2012/05/23
OBSERVER: Paul Ellen
CONDITION: Clear -> some cloud
COMMENT: The HDD backup is still running at 20:00... Waited for HDD backup to finish before observing start.
Compressor gas pressure: 1.7 MPa
Expected Observation Time: START 18:26 END 07:03
Go to the Seeing Log
Go to the Observation Log
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 37023-37071 | 49 | 325 | 20700 |
LMC | - | 0 | 0 | 0 |
SMC | - | 0 | 0 | 0 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 49 | 325 | 20700 (05h45m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
11 | 11 | 32 | 32 | 49 | 1 | 6 | 11 | 32 | 32 | 6 | 6 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
11 | 32 | 6 | 6 | 11 | 11 | 6 | 6 | 6 | 1 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 |
0 | 0 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
18:26 | waiting | Clear |
19:00 | waiting | Clear |
20:00 | observing | Clear |
21:00 | observing | Clear & some wind |
22:00 | observing | Some cloud & some wind |
23:00 | observing | Some cloud |
00:00 | observing | Some cloud |
01:00 | observing | Some cloud |
02:00 | observing | Some cloud |
03:00 | observing | Some cloud |
04:00 | observing | Some cloud |
05:00 | observing | Some cloud |
06:00 | observing | Some cloud |
07:00 | observing | Some cloud |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day